IF YOU START WITH 50N G OF IRON III OXIDE,AND 25 G AL, WHAT IS THEMAXIMUM MASS OF ALUMINUM OXIDE THAT COULD BE PRODUCED? HOW MUCH AL OXIDE WOULDNBE PRODUCED IF HE YIELD IS 93%


1
Expert's answer
2022-01-30T08:01:58-0500

Fe2O3 + 2Al = Al2O3 + 2Fe

According to the equation, n (Al2O3) = 2 n (Al) = n (Fe2O3)

M (Fe2O3) = 159.7 g/mol

M (Al) = 26.98 g/mol

M (Al2O3) = 101.96 g/mol

n (Fe2O3) = m / M = 50 / 159.7 = 0.31 mol

n (Al) = 25 / 26.98 = 0.93 mol

Fe2O3 is the limiting reactant.

n (Al2O3) = n (Fe2O3) = 0.31 mol

m (Al2O3) = n x M = 0.31 x 101.96 = 31.61 g

m (Al2O3)93% = 31.61 x 0.93 = 29.4 g